find a polynomial with leading coefficient 1 and degree 3 that has 1,1,and 3 as roots.
Answer by jcmtnez(53)

That polynomial can be expressed as:
(x-1)(x-1)(x-3) Now we have to expand this expression
(x-1)(x-1)(x-3)=(x^2-2x+1)(x-3)=x^3-2x^2+x-3x^2+6x-3=x^3-5x^2+7x-3
So the polynomial of degree three that have the solution (1,1,3) is
x^3-5x^2+7x-3
I'm glad to help you anytime